# Finite Element and Machine Learning Modeling of Autogenous Self-Healing Concrete

### v1

@@ 3.4 Comparison between CDM and CMM
Comparing the two models in Figure 11, the added functionality of halting water movement until sufficient moisture is met in the Crack Membrane Model significantly lengthens the healing time. In the beginning (during Stage 1), the Crack Membrane Model exhibits a substantially slower initial healing rate compared to the Crack Diffusion Model. During Stage 3, the Crack Membrane Model achieves healing rates comparable to the final healing rates of the Crack Diffusion Model. The Crack Membrane Model provides a more accurate representation of the underlying physical and chemical processes. Experimental validation should demonstrate these characteristic transitions and distinct healing stages. However, the Crack Membrane Model’s additional calibration requirements increase analytical complexity. Thus, the Crack Diffusion Model was chosen as the primary model for the other analyses.
@@ 3.5 Performance Comparison of Machine Learning Models
The machine learning models were trained to predict the binary healing success H (0/1) based on the crack width $\sigma$, cement availability smoothing parameter $\gamma$, and time $t$. A dataset of 1,000,000 simulation samples was generated from the CDM using parameter ranges detailed in Table 2. Other parameters were kept constant. Specifically, their values were $\alpha$ = 0.01, $\beta=\frac{\pi}{2}$, $\sigma$ = 0.0224, and $\gamma$ = 0.0316. The dataset was partitioned into training (60%), validation (20%), and test (20%) subsets for model development and evaluation. Performance metrics including accuracy, precision, recall, and f1-score [21] were computed on the test subset. The performance comparison is shown in Table 3 and Figure 13.
Parameter Lower Bound Upper Bound Step Size Number of Steps
$\sigma$ $0.001\text{ cm}$ $1\text{ cm} 0.0$$\text{ cm} 100100$
$\gamma$ $0.001\text{ cm}$ $1\text{ cm} 0.0$$\text{ cm} 100100$
[…]
Table 3: Machine Learning Model Performance Comparison
Metric KNN Naive Bayes Logistic Regression SVM Neural Network
Accuracy 0.9971 0.9481 0.9879 0.9936 0.9980
Precision 0.9972 0.9618 0.9886 0.9939 0.9980
Recall 0.9971 0.9481 0.9879 0.9936 0.9980
F1-Score 0.9972 0.9513 0.9881 0.9937 0.9980
Figure 13: Machine Learning Model Performance Comparison Bar Chart (accuracy, precision, recall, f1-score)
The neural network achieved the highest performance across all metrics (accuracy=0.9980, precision=0.9980, recall=0.9980, f1-score=0.9980) using a 3-32-32-1 architecture with ReLU activation, dropout=0.2, learning rate=0.001, batch size=32, and 100 epochs. The performances of KNN, SVM, Logistic Regression, and Naive Bayes followed, in that order. It should be noted that the neural network took the longest time to train, while the KNN algorithm was almost instantaneous. Thus, for maximal performance, the neural network is the best pick. For a combination of performance and training time, KNN is the best pick.
@@ 3.6 Limitations and Future Improvements
Future work should address these limitations through experimental validation, realistic crack geometry implementation, and incorporation of porous media flow mechanics. Additionally, the model could be extended to three-dimensional domains to capture volumetric healing effects, cyclic loading scenarios to assess long-term durability, and strength recovery calculations through post-processing of damage fields. These extensions would enhance the model’s predictive capability for real-world applications. The ML models can be further analyzed using explainable AI (XAI) to provide insights into the model’s predictions.
@@ 4 Conclusions
This study developed two coupled finite element models for simulating autogenous self-healing concrete behavior, addressing the critical need for computational tools to evaluate this promising technology for infrastructure reenforcement and sustainability. The Crack Diffusion Model (CDM) and Crack Membrane Model (CMM) successfully integrated time-dependent feedback mechanisms, logarithmic interpolation of diffusion coefficients, and damage evolution based on cement and water availability. CMM adds a binary gate mechanism to the CDM to simulate the moisture threshold buildup and rapid healing after gate activation.
The models revealed several key insights into healing behavior. Crack orientation significantly influences healing time, with diffusion distance to crack endpoints being the dominant factor rather than crack length. Healing time increases from $0^{\circ}$ to $45^{\circ}$ as water travels farther to reach crack endpoints, then remains relatively constant from $45^{\circ}$ to $135^{\circ}$ due to similar diffusion distances, with a slight decrease at $90^{\circ}$ due to shorter crack length. The relationship between crack width and healing time depends on the relative diffusivities of cracked versus intact regions, demonstrating the importance of material parameter selection.
@@
Machine learning models trained on over one million simulation data points achieved high accuracy (up to 99.8% for neural networks) in predicting healing outcomes, demonstrating the potential for rapid screening of design parameters. The k-nearest neighbors algorithm provided an optimal balance between accuracy (99.7%) and computational efficiency.
@@
The computational framework developed in this study could accelerate the adoption of self-healing concrete technologies, potentially reducing the environmental and economic costs associated with infrastructure maintenance and repair. Future work should focus on experimental calibration, three-dimensional extensions, and cyclic loading scenarios to further enhance predictive capability for real-world applications.

### v2

@@ 3.4 Comparison between CDM and CMM
Comparing the two models in Figure 13, the added functionality of halting water movement until sufficient moisture is met in the Crack Membrane Model significantly lengthens the healing time. In the beginning (during Stage 1), the Crack Membrane Model exhibits a substantially slower initial healing rate compared to the Crack Diffusion Model. During Stage 3, the Crack Membrane Model achieves healing rates comparable to the final healing rates of the Crack Diffusion Model. The Crack Membrane Model provides a more accurate representation of the underlying physical and chemical processes. Experimental validation should demonstrate these characteristic transitions and distinct healing stages.
Previous studies have shown that standard linear diffusion approaches are inadequate for cracks under viscous or inertial pressure gradients. For these cases, crack width and roughness determine the nonlinear flow of water [32]. While past models have documented pressure-based water flow through cracks, the present work applies the mechanism to self-healing concrete.
The Crack Membrane Model’s additional calibration requirements increase analytical complexity. Thus, the Crack Diffusion Model was chosen as the primary model for the other analyses.
@@ 3.5 Performance Comparison of Machine Learning Models
Figure 15: Machine Learning Model Performance Comparison Bar Chart ($R^{2}$)
The $R^{2}$ score performance comparison for the five machine learning models is shown in Figure 15. The Gradient Boosting Model (XGBoost) achieved the highest performance with an $R^{2}$ score of 0.999952, slightly higher than the Random Forest Model with an $R^{2}$ score of 0.999938. The performances of the Gaussian Process, Neural Network, and Ridge Regression Models followed, in that order. Previous studies have demonstrated that regression-based machine learning models can predict self-healing behavior in concrete with $R^{2}$ values typically exceeding $0.9$ [33]. The present model achieves an $R^{2}$ value of $0.999952$, which surpasses these prior benchmarks and demonstrates sufficient accuracy to serve as an effective surrogate for finite element simulations.
Figure 16: SHAP Analysis of XGBoost Machine Learning Model (a) Mean Absolute SHAP Value Comparison Bar Chart, (b) SHAP Feature Impact Distribution Summary Plot
SHAP was used to analyze the XGBoost model. The analysis results are shown in Figure 16. The mean absolute SHAP value comparison bar chart (Figure 16(a)) quantifies the average magnitude of each feature’s contribution, revealing that crack orientation $\beta$ exhibits the highest mean absolute impact on healing time predictions at $5.85\times 10^{5}$ seconds, followed by crack width indicator $\sigma$ at $3.88\times 10^{4}$ seconds and cement availability smoothing parameter $\gamma$ at $6.21\times 10^{3}$ seconds. This observation is consistent with experimental findings reporting that healing efficiency and rate are strongly governed by the initial damage distribution [23]. Because crack orientation $\beta$ and crack width indicator $\sigma$ exert the largest and second-largest influence on the damage distribution, respectively, it is reasonable that they also exert the largest and second-largest influence on the healing time.
The SHAP feature impact distribution summary plot (Figure 16(b)) provides detailed insights into the directional and value-dependent influence of each feature. Higher values of $\beta$ (in red) tend to increase the predicted healing time (positive SHAP values), while lower values (in blue) tend to decrease the predicted healing time (negative SHAP values). The opposite is observed for $\sigma$ and $\gamma$.
Crack orientation $\beta$ demonstrates a predominantly left-skewed SHAP value distribution with lower values sparsely distributed on the negative side, while higher values are densely concentrated on the positive side. This reinforces the notion that the healing time is diffusion-time limited, since the diffusion distance to the crack endpoint for $\beta=45^{\circ}$ to $\beta=90^{\circ}$ is approximately constant at $1$ cm.
Crack width indicator $\sigma$ exhibits an approximately symmetric range centered near zero, but shows a large cluster of negative SHAP values for higher values of $\sigma$. This relationship aligns with Figure 10 and Figure 11, where higher $\sigma$ values result in a decrease in the healing time. As $\sigma$ continues to increase, the decrease in healing time accelerates, causing a slight right-skew in the SHAP value distribution.
Cement availability smoothing parameter $\gamma$ displays a strong right-skewed SHAP value distribution. As $\gamma$ increases, the healing time decreases. This relationship is consistent with Figure 11, which shows that healing time exhibits a convex, decreasing dependence on $\gamma$.
@@ 3.6 Limitations and Future Improvements
Future work should address these limitations through experimental validation, realistic crack geometry implementation, and incorporation of porous media flow mechanics. Validation against experimental data is essential to verify simulation reliability. This validation should include direct comparison of predicted healing times and damage evolution patterns with laboratory measurements. Mesh convergence studies and comparison with analytical solutions will further establish the numerical accuracy of the finite element analysis [34]. The models could be extended to three-dimensional domains to capture volumetric healing effects, cyclic loading scenarios to assess long-term durability, and strength recovery calculations through post-processing of damage fields. Incorporation of additional healing factors, including concrete hydration age, environmental curing conditions, curing history, and evolving microstructure, would further enhance the model’s predictive capability for real-world applications. The machine learning models can be further optimized using physics informed neural networks (PINNs), incorporating a loss function that penalizes the model’s predictions if it violates the diffusion and healing processes.
@@ 4 Conclusions
This study developed two coupled finite element models for simulating autogenous self-healing concrete behavior, addressing the critical need for computational tools to evaluate this promising technology for infrastructure reenforcement and sustainability. The Crack Diffusion Model (CDM) and Crack Membrane Model (CMM) successfully integrated time-dependent feedback mechanisms, logarithmic interpolation of diffusion coefficients, and damage evolution based on cement and water availability. The models implement a Helmholtz filtering approach to model the distance over which cement clinker can travel to form crystals during the healing process. CMM adds a threshold-based binary gate mechanism to the CDM to simulate the moisture threshold buildup and rapid healing after gate activation. This study developed a machine learning framework that serves as a more computationally efficient surrogate for the finite element models.
The models revealed several key insights into healing behavior:
  • Crack orientation significantly influences healing time, with diffusion distance to crack endpoints being the dominant factor rather than crack length. Healing time increases from $0^{\circ}$ to $45^{\circ}$ as water travels farther to reach crack endpoints, then remains relatively constant from $45^{\circ}$ to $135^{\circ}$ due to similar diffusion distances, with a slight decrease at $90^{\circ}$ due to shorter crack length.
  • The relationship between crack width and healing time depends on the relative diffusivities of cracked versus intact regions. If the diffusion coefficient in cracked regions is less than that in intact regions, the healing time increases with crack width. If the diffusion coefficient in cracked regions is greater than that in intact regions, the healing time decreases with crack width.
  • A 3D surface plot revealing the intertwined influence of crack width and cement availability smoothing on healing time showed that healing time decreases as the smoothing parameter increases. Changing the smoothing parameter preserved the shape of the healing time versus crack width function, but shifted it along a convex, decreasing curve.
@@
  • Machine learning models trained on over one thousand simulation data points achieved high accuracy (up to $R^{2}=0.999952$ for the Gradient Boosting Model) in predicting healing time, demonstrating the potential for rapid screening of design parameters.
  • SHAP analysis demonstrated that crack orientation influenced the healing time the most, followed by crack width and cement availability smoothing. This demonstrates that the healing time is diffusion-time limited, since angle and width of the crack are the most important factors in determining the diffusion distance. The feature impact distribution summary plot aligned with results from the finite element analysis.
@@
Additional factors that can be modeled include concrete hydration age, environmental curing conditions, curing history, and evolving microstructure. Future work should focus on experimental calibration, three-dimensional extensions, and cyclic loading scenarios to further enhance predictive capability for real-world applications.
The computational framework developed in this study could accelerate the implementation of self-healing concrete technologies by enabling rapid parameter screening and optimization before expensive experimental trials, providing guidance for material design, and reducing the time and cost required to evaluate healing performance across diverse conditions. The machine learning surrogate models further enhance this capability by delivering near-instantaneous predictions. This combined finite element and machine learning approach reduces the risk and uncertainty associated with field implementation, thereby accelerating the adoption of self-healing concrete, which has the potential to reduce the environmental and economic costs associated with infrastructure maintenance and repair.
CRediT Author Contribution Statement
William Liu: Conceptualization, Data Curation, Formal Analysis, Investigation, Methodology, Project Administration, Resources, Software, Validation, Visualization, Writing - Original Draft, Writing - Review and Editing.
Declaration of Competing Interest
The author declares that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.